OpenStudy (anonymous):

show that √((1-x)/(1+x))=1-x+1/2(x^2)

OpenStudy (thomas5267):

\(\sqrt{\frac{1-x}{1+x}}=1-x+\frac{1}{2(x^2)}\)?